OpenStudy (anonymous):

The amount of oil used by a ship traveling at a uniform speed varies jointly with the distance and the square of the speed. The ship uses 30 barrels of oil in traveling 85 miles at 42 mi/h. How many barrels of oil are used when the ship travels 26 miles at 54mi/h? Round your answer to the nearest tenth of a barrel if necessary.

OpenStudy (blake57roger):

30=k * 85*(42)^2 30 = 149940k k = 30/149940 = 3/14994 = 1/4998 d = 26 s = 54 O = (26•54^2)/4998 ≈ 15.2

OpenStudy (anonymous):

Ok thanks! I'm glad I did it right. The question was confusing, but I came to that same conclusion. @blake57roger

OpenStudy (blake57roger):

thanks